OpenStudy (anonymous):

Compute the derivative of cos(sqrt(x)) , with respect to x from first principle.

OpenStudy (anonymous):

@experimentX @Mashy

OpenStudy (anonymous):

it would be \[-1/2 \sin x\]

OpenStudy (anonymous):

I am not 100 percent sure

OpenStudy (anonymous):

isnt the derivative of cos(sqrt(x)) = -sin(sqrt(x))/2sqrt(x)

OpenStudy (anonymous):

yea that seems right

OpenStudy (johnweldon1993):

i believe so @ton12

OpenStudy (anonymous):

y u asking if you know lol

OpenStudy (anonymous):

But Hw do u Figure out with 1st principle it is getting werid with that

OpenStudy (anonymous):

\[-1/2 x ^{-1/2}\sin \sqrt{x}\]

OpenStudy (anonymous):

i want to figure out it from 1st principle @aajugdar

OpenStudy (anonymous):

uh well differentiate root x then differentiate sin root x then multiply

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

What is first principle?

OpenStudy (anonymous):

lim h--->0 ( f(x+h) - f(x) ) / h

OpenStudy (anonymous):

oh lol okay lets see

OpenStudy (anonymous):

Lt h->0 [cos (√(x+h) - cos √x) ]/h = - 1/h *2 * sin (√x+h +√x) /2 sin (√(x+h) -√x)/2 = - (2sin √x) lt h-->0 1/h *sin (h/2 ((√x+h +√x) ) = - (2sin √x) * sin (h/2 ((√x+h +√x) ) ----------------------------------------… 2 ((√x+h +√x) * (h/2 ((√x+h +√x) ) - (2sin √x) * 1/(2* 2√x) * 1 = -1/2√x * sin √x

OpenStudy (anonymous):

try to get it lol

OpenStudy (anonymous):

http://www.meritnation.com/discuss/question/1757465

OpenStudy (anonymous):

check on the site you will understand

OpenStudy (anonymous):

thxx buddy

OpenStudy (anonymous):

you are welcome :)